OpenStudy (anonymous):

a1 = 1, an = an-1 + 2n + 3, n≥2 Find the closed formula

OpenStudy (anonymous):

Your sequence is this; 1, 8, 16, 26, 39, 56, 78, 106, ... Which is equivalent to this: 1 = 7n+1 8 = 7n+1 16= 7n+2 26= 7n+5 39= 7n+11 56= 7n+21 78= 7n+36 106=7n+57 1,1,2,5,11,21,36,57, ... This page describes how to generate the sequence: http://oeis.org/A050407 If you're doing this as part of a course, you may be expected to use a more specific method of solving the recurrence. You should probably specify the method you're supposed to be using so we can help you understand it.